3. The Frostbite Manufacturing Co. uses a standard cost system, and calculates the related variances. The following data pertain to material and labor costs for the manufacture of its only product: Refrigerant 410A canisters.

    Standard costs:                           

    Direct materials:   2 pounds @ $8/pound

    Direct labor:        1/2 hour @ $10/hour

        Actual results for June:  

    600 canisters produced

    Direct materials:   $10,000 for 1,290 pounds ($7.75/pound)

    Direct labor:          $ 3,300 for 320 hours ($10.31/hr.)

    Complete the following table. The cell values should represent total dollars, not per unit amounts. Label the variances F or U. (20 points.)

                                                                  

                                                 

AQ X AP

Price Variance

AQ X SP

Quantity Variance

SQ X SP

D. materials

D. labor
Answer #1

AP = 7.75 per pound

Actual Quantity = 1290

Stadard quantity = 600 * 2 = 1200 pounds

Stanadrd rate = 8 per pound

AQ*SP = 1290 * 8 = 10320

SQ*SP = 1200*8 = 9600

D. Labour = 600/2 = 300 * 10 = 3000

AQ*AP (a) Price Variance (A-b) AQ*SP (b) Quantity Variance ( b-c) SQ*SP ( C )
D.M 10000 320 Favorable 10320 -720 9600
D. Labour 3300 - 100 Unfavorable 3200 -200 3000